OpenStudy (anonymous):

what is x when x+8/-3=17-x/-2

OpenStudy (anonymous):

I would do this like this: (2*(x+8))/3 = 17 - x multiply by 2 on both sides 2(x+8) = 3*17-3x multiply by 3 on both sides 2x+16 = 51- 3x 5x = 51-16 move +16 to the right and -3x to the left x = 35/5 divide by 5 x = 5